OpenStudy (anonymous):

Rewrite (x²-3x+2)/√(x) as a sum using fractional exponents.

OpenStudy (debbieg):

Do you know how to write \(\Large \frac{ 1 }{ \sqrt{x} }\) with a fractional exponent?

OpenStudy (debbieg):

Write as \(\Large (x²-3x+2)\cdot\dfrac{ 1 }{ \sqrt{x} }\) where you write \(\dfrac{ 1 }{ \sqrt{x} }\) with a fractional exponent. Then distribute through, and combine the products using rules of exponents. \(\Large a^m\cdot a^n=a^{m+n}\) works whether the exponents are integers of rational numbers (fractions).